Can someone help me?!?! How can I find the dimensions of a rectangular box with a width three times the height, the length is 6 inches more than the height, and the volume is 15 cubic inches?
Thank You!!

Answer by venugopalramana(3286)

How can I find the dimensions of a rectangular box with a width=W three times the height=H, the length=L is 6 inches more than the height, and the volume=V is 15 cubic inches?
W=3*H
L=H+6
V=15=LWH=(H+6)3*H*H=3H^3+18H^2
H^3+6H^2-5=0
H+1 IS A FACTOR
-1|1....6.....0.....-5
...0...-1....-5......5
--------------------------------
...1....5.....-5....0
H^2+5H-5=0
H={-5+SQRT(25-20)}/2
=-VE. CHECK THE PROBLEM.....DATA GIVEN DOES NOT LEAD TO A FEASIBLE SOLUTION